OpenStudy (anonymous):

A fair five-sided spinner is numbered 1,3,5,7 and 9. Katrina spins it three times. Calculate the probability that the three numbers obtained are the same. Express as a fraction in its lowest terms.

OpenStudy (blockcolder):

It's fair, so the probability is (1/5)^3=1/125.

OpenStudy (anonymous):

@blockcolder s ryt

Parth (parthkohli):

I'm not sure about this but the 5th row of pascal's suggests that it is 1/32 :/

OpenStudy (anonymous):

I calculated 1/125 as well. But the marking scheme says it's 1/25. >_<

OpenStudy (blockcolder):

Let's do this differently. There are 5 ways to spin the wheel such that 3 #s are the same. There are 125 (5^3) possible results on the wheel. Thus, the probability that 3 #s are the same is 5/125=1/25. So I was wrong on my first try.

OpenStudy (anonymous):

5 ways to spin the wheel? I didn't get that.

OpenStudy (anonymous):

it supposed to be an unbiased, i think 1/125 is the corrct answer

OpenStudy (blockcolder):

I mean there are 5 possible results wherein the 3 numbers are the same: 111, 333, 555, 777, 999.

OpenStudy (anonymous):

OH! Now I get it. Thank you! :D

OpenStudy (blockcolder):

No problem. :D

OpenStudy (anonymous):

another way to look at this: makes no difference what the first spin is... so probability = 1 now what is the probability of the second spin matching the first? 1 in 5 and what is the probability of the third one matching them? 1 in 5 so it is 1 (1/5) (1/5) = 1/25